# Scaling Laws for Template-Free Detection of Environmental Phase Modulation in Gravitational-Wave Signals

### v1

@@ 5.3 Practical Use in Detection Pipelines
5.4 Implications for Current and Future Detectors
From the fitted sigmoid in Sec. 4, we obtain characteristic thresholds
@@ 5.4 Implications for Current and Future Detectors
For ground-based detectors such as Advanced LIGO and Virgo, typical binary black hole events have network $\mathrm{SNR}\sim 8$2020. At $\mathrm{SNR}\approx 10$, the condition $\Lambda\gtrsim 20$ implies that cumulative phase distortions of order $\Delta\phi_{\mathrm{env}}\gtrsim 2\,\mathrm{rad}$ are required for robust detectability. For louder events with $\mathrm{SNR}\approx 20$, the threshold relaxes to $\Delta\phi_{\mathrm{env}}\gtrsim 1\,\mathrm{rad}$.
@@
5.5 Limitations and Future Work
The present analysis employs controlled chirp families and a simplified constant-acceleration model. Real hierarchical triple systems can produce time-dependent accelerations beyond quadratic time shifts, and realistic compact-binary waveforms contain additional structure absent from the simplified model considered here.

### v2

@@ V.3 Practical Use in Detection Pipelines
V.4 Implications for Current and Future Detectors
From the fitted sigmoid in Sec. IV, we obtain characteristic thresholds
@@ V.4 Implications for Current and Future Detectors
For ground-based detectors such as Advanced LIGO and Virgo, typical binary black hole events have network $\mathrm{SNR}\sim 8$2020. At $\mathrm{SNR}\approx 10$, the condition $\Lambda\gtrsim 20$ implies that cumulative phase distortions of order $\Delta\phi\gtrsim 2\,\mathrm{rad}$ are required for robust detectability. For louder events with $\mathrm{SNR}\approx 20$, the threshold relaxes to $\Delta\phi\gtrsim 1\,\mathrm{rad}$.
@@
V.5 Limitations and Future Work
The present analysis employs controlled chirp families and a simplified constant-acceleration model in order to isolate the time-reparameterization effect under well-defined statistical conditions. Environmental acceleration acts primarily as a smooth time warp, producing a cumulative phase deformation without strongly altering the amplitude structure of the waveform. Because this deformation is determined primarily by the accumulated phase shift $\Delta\phi_{\rm env}$, the detectability scaling with $\Lambda=\Delta\phi_{\rm env}\times\mathrm{SNR}$ is expected to hold for any smoothly evolving inspiral waveform. The post-Newtonian TaylorT4 example shown in Fig. 2 illustrates that the same trajectory-level distortion arises for physically motivated gravitational-wave signals.
Real hierarchical triple systems can produce time-dependent accelerations beyond quadratic time shifts, and realistic compact-binary waveforms contain additional structure absent from the simplified model considered here.